OpenStudy (anonymous):

Evaluate the expression when m = 4. 35 – 3m A. 1 B. 13 C. 23 D. 128

OpenStudy (anonymous):

35 – 3(4) 3 * 4 = 12 35 - 12 = 23 m = 23

OpenStudy (solomonzelman):

At first you got to know that when you say "3m" you mean "3 times m" So, your expression is \(\large\color{black}{ 35-3\times m }\) . And based on the order of operations, you multiply first. So first multiply and then subtract. Your question gives an additional information though, it says that m=4, so plug in 4 instead of m (replace "m" with 4). \(\large\color{black}{ 35-3\times 4 }\) Now just multiply \(\large\color{black}{ 35-12 }\) And lastly subtract. \(\large\color{black}{ 23. }\)